# Effective Potential and Topological Photon Spheres: A Novel Approach to Black Hole Parameter Classification

### v2

@@ 1 Introduction
In the study of ultra-compact gravitational objects, the simplest and most fundamental step to distinguish between different configurations is to examine the metric function and subsequently identify the event horizon. The metric function is essentially the result of integrating equations based on diverse and complex initial conditions (such as energy conditions, geometric compatibility with general relativity equations, etc.). Any remaining weaknesses that might challenge the physics of the derived model are addressed by conjectures such as the Weak Cosmic Censorship Conjecture (WCCC), which posits the creation of an event horizon. This conjecture assumes that singularities emerging from solving Einstein’s equations must be hidden behind an event horizon. This is because if singularities were observable from the rest of spacetime, causality would be compromised, and physics would lose its predictive power. It appears that this condition is more of an attempt to cover current scientific and instrumental weaknesses and our lack of understanding of geometric spacetime, which we are actually unable to describe. It is important to note that the failure to meet the WCCC condition does not imply ”non-existence” or ”absence,” but rather the ”existence” of a spacetime that, although it has its own gravitational effects, our current knowledge is insufficient to decode it. Extensive studies are currently being conducted on the possibility and conditions for the emergence and observation of these regions. For instance, spacetime singularities formed during gravitational collapse can be observed by an external observer, and the observability of a spacetime singularity depends on the initial conditions of the collapsing matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Therefore, it is crucial to know under what conditions our model behaves like a black hole and from what range it transitions to an ultra-compact object without an event horizon (naked singularity). However, naked singularity regions may also be involved under certain conditions in the results. Hence, their extent of influence must also be determined.
@@
After constructing the initial metric function, the most significant factor influencing the location of the event horizon is the values assigned to the parameters of each theory used in constructing the metric function. In a simplified view, the permissible range of these parameters can be studied based on the roots of the metric function. However, it is clear that this method will not be applicable in determining the dominance range of a naked singularity. Therefore, to gain a better understanding of the overall space surrounding an ultra-compact object, one must seek a potential function that provides a more comprehensive response to how the gravity of this structure affects the entire surrounding space. In pursuit of such a function, we know that the nature of gravitational structures, whether in their classical or quantum forms, necessitates the existence of stable and unstable circular orbits. Moreover, in gravity, the existence of such orbits around a black hole or an ultra-compact object has been well calculated both experimentally and observationally, as well as theoretically for various black hole models.For example, in a observational study, investigators documented the presence of a distinct dark shadow, which emerged due to the gravitational deflection of light by the colossal black hole located at the nucleus of the massive elliptical galaxy M87. This phenomenon conclusively substantiates the existence of photon rings encircling the central dense object [15, 16, 17, 18, 19, 20]. Also, theoretically, Cunha et al. showed during their studies that, in spherical symmetry, standard black holes, as well as other ultra-compact objects (that can be with or without event horizon) in general relativity, can have planar circular photon orbits. In such a way that the stable type causes instability, and its unstable type could determine the shadows of the black hole [21]. Also, in a new paper it was shown in a geometrical proof that ’any spherically symmetric space-time that is asymptotically flat and has a horizon must contain a photon sphere, a spherical photon shell’ [22]. Therefore, considering the above characteristics, it seems that the idea of the photon sphere can be used for a more precise classification of the space surrounding the ultra-compact structures.
@@
In this article, we will use the ”Wei” method to classify space based on the topological photon sphere behavior for different black holes. It means that, the topological current is non-zero only at the zero point of the vector field, which determines the location of the photon sphere. Therefore, one topological charge can be considered for each photon sphere. In the full exterior region, based on a specific range of parameters where the Total Topological Charge(TTC) is always equal to -1, the structure will be considered a black hole, and conversely. Moreover, on the same basis, for a naked singularity that has a vanishing topological charge, the TTC will be 0 [24] or +1 [25]. Ultimately, if neither of the above states occurs and, on the other hand, the metric function and effective potential within the studied range are without roots, due to the violation of the condition for the existence of a photon sphere in these regions, these areas will likely represent forbidden zones.
@@ 2 The topological path to the photon sphere
once again and this time around the topological charge Q, according to the characteristic of the $\delta$ function, it can be said that the charge is non-zero only at the zero point of $\phi$. This results lead us how to use and calculate the photon sphere in the near future. Now, we are going to investigate the photon sphere and some null geodesics with static and spherical symmetry background (1), Here, we consider four-momentum and Hamiltonian condition for null geodesics,
@@ 4 Effective potential and topological photon sphere as a tool for classification
However, there are fundamental drawbacks that must be considered. Since the vector field ($\phi$) must be separable with respect to spatial coordinates, only diagonal metrics with spherical symmetry can be used. Given that most models of interest have spherical structures and the rest can often be expressed diagonally using the C-metric, this limitation can be somewhat overlooked. Another issue is that since the Poincaré plane is used for dimensional reduction, this method may not be applicable to models with fewer than four dimensions. Finally, as this is a nascent approach, there may be weaknesses that become apparent over time. Interestingly, the simplifications, computational power, and novelty of this method and mathematical model for studying symmetric potential functions have garnered such attention that the mathematical structure of the model has been extended to thermodynamics. Consequently, extensive studies based on the mathematical model of this method have been conducted on the phase transitions of black hole models [33, 34, 35, 36, 37, 38, 39, 40, 41].
@@ 5 Conclusions
Given the necessity of photon rings and photon spheres for gravitational structures in numerous studies[15, 16, 17, 18, 19, 20, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74], this paper, for the first time, adopted a novel approach by utilizing the existence and behavioral patterns of topological charges associated with photon and anti-photon spheres. We aimed to determine the parameter ranges and their impact on the behavior of ultra-compact objects with various gravitational structures.
To achieve this, we examined models with nonlinear fields or multiple horizons, often influenced by PDFM. We meticulously plotted field diagrams, zero points, total topological charges, and effective potential diagrams. Based on these results, we provided a detailed classification of the behavior and characteristics of each model in Tables 1-7.
Overall, it is essential to state that the bidirectional relationship between the photon sphere and the gravitational structure is an ideal advantage, and studying it from both perspectives can lead to several interesting outcomes:
• As an initial result, calculating topological photon spheres can complement metric function results and even serve as a suitable alternative for examining the behavior of gravitational structures. An unstable photon sphere dominates the system’s space only when an event horizon exists. Therefore, the range in which the model behaves as a black hole essentially meets the basic requirements of the WCCC. Additionally, the broader spatial coverage of the topological photon sphere study method around the gravitational structure is a key point that distinguishes it from the metric function.
• One of the main objectives of this paper is to easily classify the parameter space of the studied models based on the existence and topological location of photon and anti-photon spheres. This means we can precisely identify which parameter range of the model corresponds to a black hole structure or a naked singularity.[26]
• According to this study, if we slice the spacetime of an ultra-compact object (a naked singularity) at the location of the photon sphere with a horizontal plane, there will be at least one ring. An unstable photon sphere with a total charge of -1 or several distinct rings (stable and unstable) with a total charge of 0 or +1. Given that any photon trapped in a stable ring never leaves its orbit, practically, from this boundary towards the center of the structure, it is an area considered invisible to us. Due to the critical importance of the photon sphere in determining the shadow, this ring essentially acts as a visible boundary for us. Now, since according to the WCCC conjecture, any unpredictable region must be hidden behind the event horizon, our question is whether, hypothetically, the radius of the stable photon sphere can be considered a new hypothetical horizon for the studied spacetime? What implications does this choice have for the structure of the ultra-compact object?
@@
• Another achievement of this method is that we can adjust the system parameters to force the studied model to exhibit the desired behavior or, conversely, estimate the approximate parameter ranges of the model based on the presence of the photon sphere in a specific region of space. For example, in some gravitational studies, we need a gravitational structure that adheres to specific principles and behaviors. In such cases, where we need the model parameters to fall within a specific range while simultaneously maintaining the structure as a defined black hole, this method can play a significant role. To further clarify, in the Swampland conjecture model study, hypotheses such as the Weak Gravity Conjecture (WGC) require that the studied structure, while being overly charged, still maintains its black hole structure. In this work, we propose that for such scenarios, the topological behavior of the photon sphere can be used, and after adjusting the parameters as desired, the topological photon sphere can be examined to determine whether optimal conditions can be achieved.

### v3

@@ 1 Introduction
In the study of ultra-compact gravitational objects, the simplest and most fundamental step to distinguish between different configurations is to examine the metric function and subsequently identify the event horizon. The metric function is essentially the result of integrating equations based on diverse and complex initial conditions (such as energy conditions, geometric compatibility with general relativity equations, etc.). Any remaining weaknesses that might challenge the physics of the derived model are addressed by conjectures such as the Weak Cosmic Censorship Conjecture (WCCC), which posits the creation of an event horizon. This conjecture assumes that singularities emerging from solving Einstein’s equations must be hidden behind an event horizon. This is because if singularities were observable from the rest of spacetime, causality would be compromised, and physics would lose its predictive power. It is important to note that the failure to meet the WCCC condition does not imply ”non-existence” or ”absence,” but rather the ”existence” of a spacetime that, although it has its own gravitational effects, our current knowledge is insufficient to decode it. Extensive studies are currently being conducted on the possibility and conditions for the emergence and observation of these regions. For instance, spacetime singularities formed during gravitational collapse can be observed by an external observer, and the observability of a spacetime singularity depends on the initial conditions of the collapsing matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Therefore, it is crucial to know under what conditions our model behaves like a black hole and from what range it transitions to an ultra-compact object without an event horizon (naked singularity). However, naked singularity regions may also be involved under certain conditions in the results. Hence, their extent of influence must also be determined.
@@
After constructing the initial metric function, the most significant factor influencing the location of the event horizon is the values assigned to the parameters of each theory used in constructing the metric function. In a simplified view, the permissible range of these parameters can be studied based on the roots of the metric function. Because all necessary conditions, including the geometric equations of general relativity and energy conditions, have been integrated within the system of equations and the result has manifested as a model characterized by a defined metric function. However, it is clear that this method will not be applicable in determining the dominance range of a naked singularity. Therefore, to gain a better understanding of the overall space surrounding an ultra-compact object, one must seek a potential function that provides a more comprehensive response to how the gravity of this structure affects the entire surrounding space. In pursuit of such a function, we know that the nature of gravitational structures, whether in their classical or quantum forms, necessitates the existence of stable and unstable circular orbits. Moreover, in gravity, the existence of such orbits around a black hole or an ultra-compact object has been well calculated both experimentally and observationally, as well as theoretically for various black hole models.For example, in a observational study, investigators documented the presence of a distinct dark shadow, which emerged due to the gravitational deflection of light by the colossal black hole located at the nucleus of the massive elliptical galaxy M87. This phenomenon conclusively substantiates the existence of photon rings encircling the central dense object [15, 16, 17, 18, 19, 20]. Also, theoretically, Cunha et al. showed during their studies that, in spherical symmetry, standard black holes, as well as other ultra-compact objects (that can be with or without event horizon) in general relativity, can have planar circular photon orbits. In such a way that the stable type causes instability, and its unstable type could determine the shadows of the black hole [21]. Also, in a new paper it was shown in a geometrical proof that ’any spherically symmetric space-time that is asymptotically flat and has a horizon must contain a photon sphere, a spherical photon shell’ [22]. Therefore, considering the above characteristics, it seems that the idea of the photon sphere can be used for a more precise classification of the space surrounding the ultra-compact structures.
@@
In this article, we will use the ”Wei” method to classify space based on the topological photon sphere behavior for different black holes. It means that, the topological current is non-zero only at the zero point of the vector field, which determines the location of the photon sphere. Therefore, one topological charge can be considered for each photon sphere. In the full exterior region, based on a specific range of parameters where the Total Topological Charge(TTC) is always equal to -1, the structure will be considered a black hole, and conversely. Moreover, on the same basis, for a naked singularity that has a vanishing topological charge, the TTC will be 0 [24] or +1 [25]. Ultimately, if neither of the above states occurs , these areas will likely represent forbidden zones. Forbidden in the sense that both the metric function has no roots and the potential function shows no gravitational effect on photons. An essential point to emphasize before the end of this section is that in this article, we utilize the effective potential which the notable characteristic of this potential is that it solely depends on the geometry and structure of spacetime, without reliance on the properties of the incoming particle. We will examine the existence of photon spheres in two scenarios: structures with an event horizon (black holes) and those without an event horizon (naked singularities). In addition to identifying the locations of these spheres and demonstrating the behavioral alignment of topological charge with the geometric behavior of the potential function, we will show how the permissible range for each parameter must be defined for the structure to mathematically manifest as either a black hole or a naked singularity. It is crucial to note that our discussion is fundamentally based on the gravitational influence of the model on the motion of photons and massless particles, as well as the structure of null geodesics. Although the existence of these orbits is a necessary condition for the structure of an ultra-gravitational object, this does not imply that the obtained range necessarily holds completely physical meaning. This range will be most meaningful when interpreted alongside other physical conditions and constraints. Overall, this method can provide a comprehensive view of the behavior of ultra-gravitational structures and serve as an auxiliary equation, offering better insight into the impact of the effective components in the black hole dynamics of the model.
@@ 2 The topological path to the photon sphere
once again and this time around the topological charge Q, according to the characteristic of the $\delta$ function, it can be said that the charge is non-zero only at the zero point of $\phi$. This results lead us how to use and calculate the photon sphere in the near future.
Now, we are going to investigate the photon sphere and some null geodesics with static and spherical symmetry background (1),the Lagrangian takes the form [23]:
$\mathcal{L}={x^{\cdot}}^{\nu}{x^{\cdot}}^{\mu}g_{\mu\nu}={t^{\cdot}}^{2}f\!\left(r\right)-\frac{{r^{\cdot}}^{2}}{g\!\left(r\right)}-{\varphi^{\cdot}}^{2}h\!\left(r\right),$
where the dot denotes a derivative with respect to an affine parameter. Now we could derive the four-momentum from this Lagrangian, we have:
$p_{\mu}=\frac{\partial}{\partial{x^{\cdot}}^{\mu}}\mathcal{L}={x^{\cdot}}^{\nu}g_{\mu\nu},$
or more precisley
$\begin{split}&p_{t}=t^{\cdot}f\!\left(r\right)\equiv E,\\ &p_{\varphi}=-\varphi^{\cdot}h\!\left(r\right)\equiv-L,\\ &p_{r}=-\frac{r^{\cdot}}{g\!\left(r\right)}\\ .\end{split}$
Where E and L respectively interpreted as the energy and angular momentum. As is evident from the Lagrange equation above, this function is independent of t and $\phi$, so $p_{t}$ and $p_{\phi}$ will be the two integrals of motion, namely:
$\begin{split}&\varphi^{\cdot}=\frac{L}{h\!\left(r\right)}\\ &t^{\cdot}=\frac{E}{f\!\left(r\right)}.\\ \end{split}$
Now for the general form of Hamiltonian we have:
$\Pi={x^{\cdot}}^{\mu}p_{\mu}-\mathcal{L}=\Delta,$
where $\Delta=1,0$ for time-like and null geodesics. Here, we consider four-momentum and Hamiltonian condition for null geodesics,
@@ 3 Explanation of topological charges based on properties of effective potential
Before delving into further details, perhaps the most important question to address is: What insights can the study of different photon sphere structures provide regarding the nature of black holes? It is important to state that the stability/instability of photon orbits around black holes plays a crucial role in shaping the observed black hole shadow, and can also provide insights into the presence of naked singularities. The shadow of a black hole is essentially the silhouette created by the event horizon against the backdrop of light from the accretion disk or other sources. The shape and size of this shadow are influenced by the photon orbits, which are paths that light can take around the black hole. The EHT’s observations of M87* and Sgr A* provide invaluable data on black hole shadows. These observations confirm the predictions of stable orbits affecting the shape and size of the shadow. The amount of light that visibly circles the black hole and contributes to its shadow can show insights into the structure surrounding it, potentially revealing the presence of an accretion disk, which influences photon paths [33]. For a non-rotating black hole, the shadow is nearly circular. However, for a rotating black hole, the shadow becomes distorted due to frame-dragging effects, where space-time itself is twisted by the black hole’s rotation [34]. The size of the shadow is determined by the radius of the photon sphere. A larger photon sphere results in a larger shadow. If photon orbits were stable, the shadow would be more sharply defined, as photons would orbit the black hole for longer periods before either escaping or being captured [35]. In reality, the unstable orbits, leading to a more diffuse shadow boundary. This instability causes the photons to either spiral into the black hole or escape, contributing to the characteristic ”fuzzy” edge of the shadow[36]. But the story about the naked singularity is completely different. Naked singularities, if they exist, would present a very different observational signature compared to black holes The lack of an event horizon means that photons could potentially orbit much closer to the singularity, leading to more extreme bending of light and potentially observable differences in the light patterns. This would lead to unique gravitational lensing effects, where light from background stars or other objects would be bent in unusual ways. In these cases, instead of well-defined shadows, one might expect unique visual signatures - such as anomalously bright phenomena formed by gravitational lensing effects or different patterns of light emissions due to their unstable nature [37].
@@ 4 Effective potential and topological photon sphere as a tool for classification
However, there are fundamental drawbacks that must be considered. Since the vector field ($\phi$) must be separable with respect to spatial coordinates, only diagonal metrics with spherical symmetry can be used. Given that most models of interest have spherical structures and the rest can often be expressed diagonally using the C-metric, this limitation can be somewhat overlooked. Another issue is that since the Poincaré plane is used for dimensional reduction, this method may not be applicable to models with fewer than four dimensions. Finally, as this is a nascent approach, there may be weaknesses that become apparent over time. Interestingly, the simplifications, computational power, and novelty of this method and mathematical model for studying symmetric potential functions have garnered such attention that the mathematical structure of the model has been extended to thermodynamics. Consequently, extensive studies based on the mathematical model of this method have been conducted on the phase transitions of black hole models [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55].
@@ 5 Conclusions
Given the necessity of photon rings and photon spheres for gravitational structures in numerous studies [15, 16, 17, 18, 19, 20, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88] this article presents a new approach for determining parameter ranges in ultra-compact gravitational objects using topological charges and behavioral patterns associated with photon spheres. To achieve this, we examined models often influenced by PDFM and, in one case, a model with multiple horizons. We meticulously plotted field diagrams, zero points, total topological charges, and effective potential diagrams. Based on these results, we provided a detailed classification of the behavior and characteristics of each model in terms of gravitational influence on photon motion, as shown in Tables 1-7. To better understand the obtained results, it is useful to compare the results presented in the tables before summarizing the overall findings. As can be seen from the comparison of Tables 1 and 2, in the PDFM black hole, the parameter $\Lambda$ only causes the model to take the form of a black hole within the range of 0 to 2. However, the addition of charge and the AdS radius completely alters this range, such that the model no longer exhibits the form of a naked singularity within the studied values. This behavioral difference, which arises solely from the addition of electric fields to the model, can be quite intriguing. On the other hand, in the Euler-Heisenberg model, the results appear to be clearer. As shown in Table 3, within a wide range of negative ”a” values, the structure takes the form of a naked singularity, and only within a small range of negative ”a” values does the structure maintain the form of a black hole. Adding PDFM to the model in this case significantly reduces the range of naked singularity and allows more negative values to keep the model in the form of a black hole. We previously mentioned that negative ”a” values could indicate the presence of repulsive interactions that counteract the usual attractive gravitational effects. This means that if the model reaches an extremal or super-extremal state under certain conditions, it would provide suitable conditions for the formation of the Weak Gravity Conjecture (WGC). It is important to emphasize that although the obtained ranges may not all have physical significance, and to have a fully physical range, they must be interpreted alongside conditions such as energy conditions. However, the mutual influence of parameters and the regions where photon spheres appear can clearly serve as a significant criterion for studying the model’s behavior. This reciprocal relationship between the photon sphere and the gravitational structure is an ideal advantage, and studying it from both perspectives can lead to several interesting outcomes:
• As an initial result, calculating the location of photon spheres, interpreting their topological charge, and determining the effective gravitational range on photon motion (black hole and naked singularity) can serve as a new approach, motivating further study of various models based on this method. For instance, in this study, we observed that an unstable photon sphere, equivalent to the presence of an unstable maximum in the potential function, only appears in the spacetime influenced by the model when an event horizon exists. Therefore, based on these results and similar experiences [26, 89, 90], it can be stated that one characteristic of black holes is having an unstable photon sphere. Additionally, the broader spatial coverage of the topological photon sphere study method around the gravitational structure is a key point that distinguishes it from the metric function, meaning that it can be used to determine the gravitational influence range (in the form of black holes + naked singularities) on photon motion.
• One of the main objectives of this article is to classify the parameter space of the studied models based on the existence and topological position of photon and anti-photon spheres. This means that we can precisely determine which parameter range of the model corresponds to a black hole structure or a naked singularity. For instance, we observed that when the structure is in the form of a naked singularity, one (or more) stable photon spheres always appear alongside an unstable photon sphere in the studied spacetime. In the effective potential representation, this manifests as the appearance of a minimum alongside a maximum. Therefore, based on the results of this study and similar studies [26, 89, 90], it can be concluded that if a stable photon sphere is observed within the studied parameter range, this may indicate the possibility of the model existing in the form of a singularity. It is important to emphasize that the meaning of all the concepts discussed so far is not that the obtained ranges necessarily and entirely hold physical meaning. These ranges will be most meaningful when interpreted alongside other pre-existing physical conditions and constraints. Overall, this method can provide a comprehensive view of the behavior of ultra-gravitational structures and can serve as an auxiliary equation, offering better insight into the impact of the effective components in the black hole dynamics of the model.
@@
• Another achievement of this method is that we can adjust the system parameters to force the studied model to exhibit the desired behavior or, conversely, estimate the approximate parameter ranges of the model based on the presence of the photon sphere in a specific region of space. For example, in some gravitational studies, we need a gravitational structure that adheres to specific principles and behaviors. In such cases, where we need the model parameters to fall within a specific range while simultaneously maintaining the structure as a defined black hole, this method can play a significant role. To further clarify, in the Swampland conjecture model study, hypotheses such as the Weak Gravity Conjecture (WGC) require that the studied structure, while being overly charged, still maintains its black hole structure. In this work, we propose that for such scenarios, the topological behavior of the photon sphere can be used, and after adjusting the parameters as desired, the topological photon sphere can be examined to determine whether optimal conditions can be achieved.[89, 90, 91, 92]